# Stellar intensity interferometry in the photon-counting regime

### v1

@@ II.2.1 Alternative derivations of the SNR
However, this theoretical function will be convolved by the electronic response of the detection chain. In this process, the integral is conserved. This is the coherence time, which is only related to the optical filter:
[…]
@@
The problem is now the following: we have a Gaussian of amplitude $C\ll 1$, rms width $\tau_{\mathrm{el}}$, with an offset of 1, a time sampling $t_{\mathrm{bin}}$ and a Gaussian noise of rms amplitude $\sigma_{\mathrm{noise}}$. The sought-after SNR is given by the statistical uncertainty $\sigma_{C}$ in the determination of $C$: $S\!N\!R=C/\sigma_{C}$ (or, equivalently, in the determination of the integral).
Integral method – Let us discuss a first method: the direct numerical integration of the data and the subsequent extraction of the coherence time. It has the advantage of not relying on any prior knowledge of the response function of the detectors. The drawback is that it does not lead to the best SNR. Indeed, the integral will give $\tau_{\mathrm{c}}=1/\Delta\nu$ on average with a statistical uncertainty that is directly the uncertainty $\sigma_{\mathrm{int}}$ of the integral of the noise (here 0), given the number of points in the distribution, $n=T_{\mathrm{w}}/t_{\mathrm{bin}}$, where $T_{\mathrm{w}}$ is the window of integration: $\sigma_{\mathrm{int}}=\sigma_{\mathrm{noise}}\sqrt{n}t_{\mathrm{bin}}$. This gives an SNR
@@
This is slightly lower than the SNR of Eq. (8), as $1/\sqrt{2\sqrt{\pi}}\approx 0.53$ (we recall that we have removed the factor $|V|^{2}/2$). Moreover, this method suffers from the same caveats as the one-parameter fit, as discussed in detail in Sec. II.3.3.
@@
where $SS$ is the sum of squared residuals, $DF$ is the number of degrees of freedom, equal to the number of data points minus the number of fitting parameters, and $\mathrm{cov}(j,j)$ is the diagonal element of the covariance matrix corresponding to the fitting parameter $j$. To go further analytically, we will now consider that there is only one fitting parameter, $C$. Then,
@@
[…]
@@ II.2.2 Numerical investigation
with $y_{0}=1$, $t_{0}=0$, $\tau_{\mathrm{el}}=1$, $C=10^{-2}$ (corresponding to $\tau_{\mathrm{c}}\simeq 0.025$), and $X_{i}$ a normally-distributed random variable of mean 0 and standard deviation $\sigma_{\mathrm{noise}}$. Moreover, we consider a sampling with $t_{\mathrm{bin}}=0.02$ and a total time window $[-T_{\mathrm{w}}/2,\;\;T_{\mathrm{w}}/2]$ with $T_{\mathrm{w}}=20$ ($10^{3}$ data points). We vary $\sigma_{\mathrm{noise}}$ for a fixed $t_{\mathrm{bin}}$. Therefore, we use Eqs. (13) and (10) to change Eq. (22) to:
@@
Similarly, we can verify the result of the integral procedure, by verifying the Eq. (14), changed (using Eqs. 13 and 10) to
@@
Figure 3 shows such comparisons, along with the numerical results with two, three and four fitting parameters. We can check that the analytical results are valid and that the one-parameter fitting procedure is the most efficient. Using 4 fitting parameters, i.e.$y_{0}$, $t_{0}$, $\tau_{\mathrm{el}}$ and $C$ (Eq. 23), is clearly bad: the SNR degrades at very high SNR and the fitting procedure is less robust at low SNR, leading to a few outlier points, for which the fit did not converge properly. Interestingly, using three ($C$, $\tau_{\mathrm{el}}$ and $t_{0}$) or only two ($C$ and $\tau_{\mathrm{el}}$) fitting parameters does not make any difference. However, it is significantly better to use only one fitting parameter ($C$). For example, for $\sigma_{\mathrm{noise}}\simeq 5\times 10^{-3}$, the SNR is 15 with two or three fitting parameters and 20 with only one (and 8 for the integral method).
II.3 Experimental investigation
II.3.1 Avoiding spurious correlations
@@
  • Finally, we have observed some spurious oscillatory correlations appearing as small oscillations centered around the bunching peak, i.e. at the optical zero delay, with a range of a few tens of ns. Their amplitude decreases when the SPADs are in RF-shilded boxes and put farther apart. We interpret this phenomenon as an electronic/radiative cross-talk: when one SPAD detects a photon, its circuit emits a radiation, which very slightly modifies the detection probability of the second SPAD. Typically, and with our shielding boxes, putting the SPADs $\sim 2-3$ m apart is sufficient to make those oscillations not visible. But this can be a problem if the available space is limited (e.g. at a telescope).
Figure 4: (a) Scheme of the setup to measure $g^{(2)}(\tau)$ without spurious correlations. At the entrance of the fibered splitter, we inject light coming from a white-light source, which has been subsequently injected in a single-mode fiber (to ensure maximum spatial coherence), spectrally filtered (with $\lambda=780$ nm and $\Delta\lambda=1$ nm), and polarized (not shown). The 10-m multimode fibers (MMF) between the splitter and the SPADs ensures that the afterglow peaks are well separated from the bunching peak, and the 40-m coaxial cable on one channel ensures that the bunching peak is far away from the TDC cross-talk. Finally, the two SPADs are separated by $\sim 3$ m to avoid radiative cross-talk. (b) Corresponding $g^{(2)}(\tau)$ measurement, with an integration time of 46 h and count rates of $1.6\times 10^{6}$ counts per second per detector. The afterglow peaks reach 1.02 in amplitude. The range $\tau<-150$ nm is used to normalize the $g^{(2)}(\tau)$ function to one.
@@ II.3.2 SNR on an artificial star
Figure 5: Signal-to-noise ratio measured in the lab with an artificial star, as a function of the number of coincidences in the correlation histogram far from the bunching peak. The dashed line is the expected $S\!N\!R_{\mathrm{fit}}=\tau_{\mathrm{c}}\sqrt{N_{\mathrm{c}}/(2\sqrt{\pi}\tau_{\mathrm{el}}t_{\mathrm{bin}})}$ with $t_{\mathrm{bin}}=10$ ps. Inset: Bunching peak at the maximum integration time. The binning has been changed to $t_{\mathrm{bin}}=60$ ps for graphical improvement.
II.3.3 Practical considerations
@@
The proper method would be to use a pulsed laser to characterize the instrument response function with very high accuracy. We have done so, but we have observed, even in laboratory conditions, very small differences (on the order of 50 ps for both the position and width) between the bunching results and the pulsed-laser characterization. This is enough to make the one-parameter fit fail. It is known that the SPAD response depends on the wavelength and on the count rate, and we have taken care to have the same parameters for both acquisitions. However, it might also depend on the temperature or other less controlled parameters at the time scale of the long acquisition time. Thermal expansion in fibers and cables also influences the delays.
@@
$S\!N\!R\approx 0.44\frac{1}{2}\alpha N_{\mathrm{ph}}(\lambda)C|V(r)|^{2}\sqrt{\frac{T_{\mathrm{obs}}}{\tau_{\mathrm{el}}}}=0.44\frac{|V(r)|^{2}}{2}\frac{F}{\Delta\nu}\sqrt{\frac{T_{\mathrm{obs}}}{\tau_{\mathrm{el}}}}\;,$ (27)
where the numerical prefactor is $\approx 0.82/\sqrt{2\sqrt{\pi}}$ (we have let the $1/2$ to show that it is for unpolarized light; the $1/2$ disappears for polarized light). Note also that the factor $|V|^{2}/(2\Delta\nu)$ is simply the integral $I$ of the bunching peak. Then the equation
@@ II.3.4 On-sky validation of the SNR formula
II.3.5 Systematic error
The SNR discussed so far is only related to the statistical uncertainty of the fit, and thus does not include any systematic error. Even in the absence of any spurious correlation, there is one source of systematic error, which is that the bunching peak is not really Gaussian. The exact instrument response function (IRF) can be well characterized with a pulsed laser as input light. With our SPAD, there is a very clear deviation from a Gaussian. Fitting the IRF by a Gaussian leads to an underestimation of the integral by a few percent. However, even in our most precise bunching measurements, up to $S\!N\!R\sim 80$ [45], we have not observed any deviation from a Gaussian. We interpret this as being due to the small drifts during the acquisition, which makes the bunching peak more Gaussian than it should be.
Note that the absolute value of the coherence time is not really what matters for astronomy; it is instead its evolution as a function of the baseline. Therefore, this systematic error is mitigated by a careful calibration of the zero-baseline correlation that uses the same fitting procedure and the same couple of detectors as for the spatial correlation measurement.
II.4 On-going developments: wavelength multiplexing with high-time resolution detectors
@@
The idea is as follows: Since it is necessary to spectrally filter the light (see Sec. II.1), most of the stellar light is not used. Instead of throwing it away, one can use it for independent $g^{(2)}$ measurements at neighboring wavelengths. If the $N$ different measurements contain the same physical information, one can simply average them, and win a factor $\sqrt{N}$ on the SNR. The possibility of using wavelength multiplexing with a large number of channels, yielding a significant increase in sensitivity, is highly dependent on the availability of high-time-resolution arrays of detectors, such as SPAD arrays [60] or SNSPD (superconducting nanowire single-photon detectors) arrays [61, 62, 63], which have only recently been available, sparking the revival of the field.
@@
The design of the spectral dispersion instrument is a delicate task and discussing it goes much beyond the scope of this article. Let us briefly mention that it has to have an excellent throughput, otherwise one already needs many channels only to break even. Moreover, one needs to be careful when using gratings because they add temporal dispersion [64]. Furthermore, when computing the resulting expected SNR, one has to pay attention to remaining in the validity range of the equation. In particular, if the light is dispersed over a large number of spectral channels, the optical bandwidth may become very small for each channel, and consequently the coherence time becomes large. With timing resolution on the order of $\sim 20$ ps or even smaller [65, 66], the condition $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$ may no longer be fulfilled. It is thus interesting to see how the SNR scaling evolves in that regime. Another critical point is that the spectra we correlate should be the same at the two (or more) telescopes, which becomes more challenging as the spectral channels get narrower. We examine these two points in the following.
II.4.1 Beyond the hypothesis $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$
Figure 6: Numerical simulations of the signal-to-noise ratio as a function of $\tau_{\mathrm{el}}$ and $\tau_{\mathrm{c}}$ in different situations. (a) $\tau_{\mathrm{el}}$ is varied and $\tau_{\mathrm{c}}$ is fixed. This corresponds to improving the resolution of the detector (towards the left). The SNR increase as $1/\sqrt{\tau_{\mathrm{el}}}$ as expected until it saturates when $\tau_{\mathrm{el}}\sim\tau_{\mathrm{c}}$. The dashed line is $S\!N\!R=0.44\tau_{\mathrm{c}}/(\sigma_{\mathrm{noise}}\sqrt{\tau_{\mathrm{el}}t_{\mathrm{bin}}})$, with $\tau_{\mathrm{c}}=1$, $t_{\mathrm{bin}}=2\times 10^{-3}$ and $\sigma_{\mathrm{noise}}=10^{-2}$. (b) The detector performance is fixed ($\tau_{\mathrm{el}}=1$) and we vary the optical bandwidth $\Delta\nu$. Although the $x$-axis is the same as in panel (a) in virtue of $\tau_{\mathrm{c}}=1/\Delta\nu$, the result is different because the number of detected photons increases with $\Delta\nu$. In the limit $\Delta\nu\,\tau_{\mathrm{el}}\gg 1$, we recover the well-known result that the SNR is independent of the optical bandwidth. The dashed line is the same equation as in (a) except that now $\sigma_{\mathrm{noise}}=10^{-2}/\Delta\nu$. In the opposite regime, the SNR increases as $\sqrt{\Delta\nu}$. (c) We still consider $\tau_{\mathrm{el}}=1$ and a fixed total wavelength range $\Delta\lambda$ corresponding to $100/\tau_{\mathrm{el}}$ in frequency, divided into $N$ independent channels. The computation is the same as in (b) but the SNR is multiplied by $\sqrt{N}$. However, above $N\sim 100$, we enter the regime $\Delta\nu\tau_{\mathrm{el}}<1$ and the SNR saturates.
@@ II.4.1 Beyond the hypothesis $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$
We consider a Gaussian $g^{(2)}(\tau)$ function  [^7] We take a Gaussian instead of a sinc2 (Eq. 11) to avoid inducing a systematic error in the fitting procedure. In a true experiment, the fit function would have to be adapted. of amplitude one and area $\tau_{\mathrm{c}}$, convolved with another Gaussian of area one and rms width $\tau_{\mathrm{el}}$. We add noise as previously and determine the SNR from a 3-parameter Gaussian fit. We report the results in Fig. 6 for three different situations. First, the coherence time is fixed, and we vary $\tau_{\mathrm{el}}$ [Fig. 6(a)]. As expected, it shows that the SNR increases when the timing resolution is improved ($\tau_{\mathrm{el}}$ decreases), but the SNR saturates when $\tau_{\mathrm{el}}$ reaches $\tau_{\mathrm{c}}$. Second, we fix $\tau_{\mathrm{el}}$, and we vary $\tau_{\mathrm{c}}$ by varying the spectral bandwidth $\Delta\nu$ [Fig. 6(b)]. In that case, the number of detected photons increases with $\Delta\nu$, but since $\tau_{\mathrm{c}}$ decreases, in the limit $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$, the SNR becomes independent of the bandwidth, as already discussed in Sec. II.1. However, in the opposite regime, the (lower) SNR increases as $\sqrt{\Delta\nu}$. This is because the photon flux increases linearly with $\Delta\nu$ and the width of the measured bunching peak is reduced linearly with $\Delta\nu$, as it is fully resolved. Since the fit uncertainty is roughly proportional to the square root of the number of points in the bunching peak, we finally obtain a $\sqrt{\Delta\nu}$ dependence in this regime. Finally, we can also fix the total wavelength range and divide it into $N$ spectral channels, and study how the SNR evolves with $N$ [Fig. 6(c)]. We expect a $\sqrt{N}$ enhancement in the regime $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$, but when $\tau_{\mathrm{c}}>\tau_{\mathrm{el}}$, the SNR saturates. This happens when the number of channels is on the order of the ratio between the total optical bandwidth and the electronic bandwidth. Beyond that point, to continue improving the SNR, one has to increase the total optical bandwidth.
@@
This limitation is implicitly taken into account in the proposal [30], for example, since they consider individual bandwidths of $\delta\lambda\approx 0.1$ nm at $\lambda\approx 500$ nm with $N\approx 2000$ channels. This corresponds to $\tau_{\mathrm{c}}\approx 8$ ps, very close to the considered timing resolution $\tau_{\mathrm{el}}\approx 10$ ps. But the total wavelength range is then 200 nm, which surely makes the interpretation of the measured visibility a challenge.
@@
Finally, note that there is also a fundamental limit in simultaneously determining the arrival time and frequency (via the spectral dispersion) of a photon [68]: $\Delta\nu\Delta t\geq 1/4\pi$ (Heisenberg or Fourier dispersion relation, where here $\Delta\nu$ and $\Delta t$ are standard deviations). This means that the passage through the dispersive optics induces a supplementary timing uncertainty. The order of magnitude is the same as the coherence time 888 The exact numerical factors depend on the definitions and spectral shape. For instance, let us consider a Gaussian spectrum of rms width $\Delta\nu$ for simplicity. Then the rms width of the $g^{(2)}(\tau)$ function is ($2\pi\sqrt{2}\Delta\nu)^{-1}$ [37] and the Fourier-limited timing uncertainty $\Delta t=(4\pi\Delta\nu)^{-1}$. The two differ by a factor $\sqrt{2}$.. Therefore, it gives another reason to say that it is useless to improve the timing resolution of detectors below $\tau_{\mathrm{c}}$.
II.4.2 Effect of mismatched spectra

### v2

@@ 2.2.1 Alternative derivations of the SNR
However, this theoretical function will be convolved by the electronic response of the detection chain. In this process, the integral of $g^{(2)}(\tau)-1$ is conserved. This integral of the bunching peak is the coherence time, which is only related to the optical filter :
$\tau_{\mathrm{c}}=\int|g^{(1)}(\tau)|^{2}d\tau=\int|s(\nu)|^{2}d\nu=\frac{1}{\Delta\nu}\,.$ (13)
@@
The problem is now the following : we have a Gaussian of amplitude $C\ll 1$, rms width $\tau_{\mathrm{el}}$, with an offset of 1, a time sampling $t_{\mathrm{bin}}$ and a Gaussian noise of rms amplitude $\sigma_{\mathrm{noise}}$. Remember that $C$ is given by Eq. (14) multiplied by the set-aside factor $|V|^{2}/2$, which is the quantity we are interested in. Therefore, the sought-after SNR is given by the statistical uncertainty $\sigma_{C}$ in the determination of $C$ : $S\!N\!R=C/\sigma_{C}$ (or, equivalently, in the determination of the integral).
Integral method – Let us discuss a first method : the direct numerical integration of the data to determine the bunching-peak area. It has the advantage of not relying on any prior knowledge of the response function of the detectors. The drawback is that it does not lead to the best SNR. Indeed, the integral will give $\tau_{\mathrm{c}}=1/\Delta\nu$ on average with a statistical uncertainty that is directly the uncertainty $\sigma_{\mathrm{int}}$ of the integral of the noise (here 0), given the number of points in the distribution, $n=T_{\mathrm{w}}/t_{\mathrm{bin}}$, where $T_{\mathrm{w}}$ is the window of integration : $\sigma_{\mathrm{int}}=\sigma_{\mathrm{noise}}\sqrt{n}t_{\mathrm{bin}}$. This gives an SNR
@@
This is slightly lower than the SNR of Eq. (9), as $1/\sqrt{2\sqrt{\pi}}\approx 0.53$ (we recall that we have removed the factor $|V|^{2}/2$). Moreover, this method suffers from the same caveats as the one-parameter fit, as discussed in detail in Sec. 2.3.3.
@@
where $SS$ is the sum of squared residuals (i.e. the sum of the squared difference between the data points and the fit), $DF$ is the number of degrees of freedom, equal to the number of data points minus the number of fitting parameters, and $\mathrm{cov}(j,j)$ is the diagonal element of the covariance matrix corresponding to the fitting parameter $j$. To go further analytically, we will now consider that there is only one fitting parameter, $C$. Then,
@@
$\sum_{i}e^{-x_{i}^{2}/\tau_{\mathrm{el}}^{2}}\simeq\frac{1}{t_{\mathrm{bin}}}\int e^{-t_{i}^{2}/\tau_{\mathrm{el}}^{2}}dt=\sqrt{\pi}\frac{\tau_{\mathrm{el}}}{t_{\mathrm{bin}}}\;.$ (21)
@@ 2.2.2 Numerical investigation
with $y_{0}=1$, $t_{0}=0$, $\tau_{\mathrm{el}}=1$, $C=10^{-2}$ (corresponding to $\tau_{\mathrm{c}}\simeq 0.025$), and $X_{i}$ a normally-distributed random variable of mean 0 and standard deviation $\sigma_{\mathrm{noise}}$. Moreover, we consider a sampling with $t_{\mathrm{bin}}=0.02$ and a total time window $T_{\mathrm{w}}/2,\;\;T_{\mathrm{w}}/2]$ with $T_{\mathrm{w}}=20$ ($10^{3}$ data points). We vary $\sigma_{\mathrm{noise}}$ for a fixed $t_{\mathrm{bin}}$. Therefore, we use Eqs. (14) and (11) to change Eq. (23) to :
@@
Similarly, we can verify the result of the integral procedure, by verifying the Eq. (15), changed (using Eqs. 14 and 11) to
@@
[…]
2.3 Experimental investigation
We now turn to the application of the previous results to real-world setups.
2.3.1 Avoiding spurious correlations
@@
  • Finally, we have observed some spurious oscillatory correlations appearing as small oscillations centered around the bunching peak, i.e. at the optical zero delay, with a range of a few tens of ns. Their amplitude decreases when the SPADs are in RF-shielded boxes and put farther apart. We interpret this phenomenon as an electronic/radiative cross-talk : when one SPAD detects a photon, its circuit emits a radiation, which very slightly modifies the detection probability of the second SPAD. Typically, and with our shielding boxes, putting the SPADs $\sim 2-3$ m apart is sufficient to make those oscillations not visible. But this can be a problem if the available space is limited (e.g. at a telescope).
Figure 4: (a) Scheme of the setup to measure $g^{(2)}(\tau)$ without spurious correlations. At the entrance of the fibered splitter, we inject light coming from a white-light source, which has been subsequently injected in a single-mode fiber (to ensure maximum spatial coherence), spectrally filtered (with $\lambda=780$ nm and $\Delta\lambda=1$ nm), and polarized (not shown here, see e.g. [40]). The 10-m multimode fibers (MMF) between the splitter and the SPADs ensures that the afterglow peaks are well separated from the bunching peak, and the 40-m coaxial cable on one channel ensures that the bunching peak is far away from the TDC cross-talk. Finally, the two SPADs are separated by $\sim 3$ m to avoid radiative cross-talk. (b) Corresponding $g^{(2)}(\tau)$ measurement, with an integration time of 46 h and count rates of $1.6\times 10^{6}$ counts per second per detector. The afterglow peaks reach 1.02 in amplitude. The range $\tau<-150$ nm is used to normalize the $g^{(2)}(\tau)$ function to one.
@@ 2.3.2 SNR on an artificial star
Figure 5: Signal-to-noise ratio measured in the lab with an artificial star, as a function of the number of coincidences in the correlation histogram far from the bunching peak. The dashed line is the expected $S\!N\!R_{\mathrm{fit}}=\tau_{\mathrm{c}}\sqrt{N_{\mathrm{c}}/(2\sqrt{\pi}\tau_{\mathrm{el}}t_{\mathrm{bin}})}$ with $t_{\mathrm{bin}}=10$ ps. Inset : Bunching peak at the maximum integration time. The binning has been changed to $t_{\mathrm{bin}}=60$ ps for graphical improvement. The goodness of fit is $R^{2}=0.93$.
2.3.3 Practical considerations
@@
The proper method would be to use a pulsed laser (with pulse duration much shorter than $\tau_{\mathrm{el}}$) to characterize the instrument response function with very high accuracy. We have done so, but we have observed, even in laboratory conditions, very small differences (on the order of 50 ps for both the position and width) between the bunching results and the pulsed-laser characterization. This is enough to make the one-parameter fit fail. It is known that the SPAD response depends on the wavelength and on the count rate, and we have taken care to have the same parameters for both acquisitions. However, it might also depend on the temperature or other less controlled parameters at the time scale of the long acquisition time. Thermal expansion in fibers and cables also influences the delays.
@@
$S\!N\!R\approx 0.44\frac{1}{2}\alpha N_{\mathrm{ph}}(\lambda)A|V(r)|^{2}\sqrt{\frac{T_{\mathrm{obs}}}{\tau_{\mathrm{el}}}}=0.44\frac{|V(r)|^{2}}{2}\frac{F}{\Delta\nu}\sqrt{\frac{T_{\mathrm{obs}}}{\tau_{\mathrm{el}}}}\;,$ (28)
where the numerical prefactor is $\approx 0.82/\sqrt{2\sqrt{\pi}}$ (we have left the $1/2$ to show that it is for unpolarized light ; the $1/2$ disappears for polarized light). Note also that the factor $|V|^{2}/(2\Delta\nu)$ is simply the integral $I$ of the bunching peak. Then the equation
@@ 2.3.4 On-sky validation of the SNR formula
2.3.5 Systematic error
The SNR discussed so far is only related to the statistical uncertainty of the fit, and thus does not include any systematic error. Even in the absence of any spurious correlation, there is one source of systematic error, which is that the bunching peak is not really Gaussian, but typically contains small but large tails, whose integral is not always negligible. These tails are usually hidden in the noise and thus not counted in the bunching peak area, which is therefore underestimated. The SNR formula based on the Gaussian response hypothesis is then inaccurate.
The exact instrument response function (IRF) can be experimentally characterized with a pulsed laser as input light. With our SPADs, there is a very clear deviation from a Gaussian. Fitting the IRF by a Gaussian leads to an underestimation of the integral by a few percent. However, even in our most precise bunching measurements (using a continuous source and integrating for several days), up to $S\!N\!R\sim 80$ [48], we have not observed any deviation from a Gaussian. We interpret this as being due to the small drifts during the acquisition, which makes the bunching peak more Gaussian than it should be.
Note that the absolute value of the bunching peak area is not really what matters for astronomy ; it is instead its evolution as a function of the baseline. Therefore, this systematic error is mitigated by a careful calibration of the zero-baseline correlation that uses the same fitting procedure and the same couple of detectors as for the spatial correlation measurement.
2.4 On-going developments : wavelength multiplexing with high-time resolution detectors
@@
The idea is as follows : Since it is necessary to spectrally filter the light (see Sec. 2.1), most of the stellar light is not used. Instead of throwing it away, one can use it for independent $g^{(2)}$ measurements at neighboring wavelengths. If the $N$ different measurements contain the same physical information, one can simply average them, and win a factor $\sqrt{N}$ on the SNR. The possibility of using wavelength multiplexing with a large number of channels, yielding a significant increase in sensitivity, is highly dependent on the availability of high-time-resolution arrays of detectors, such as SPAD arrays [58] or SNSPD (superconducting nanowire single-photon detectors) arrays [59, 60, 61], which have only recently been available, sparking the revival of the field.
@@
The design of the spectral dispersion instrument is a delicate task and discussing it goes much beyond the scope of this article. Let us briefly mention that it has to have an excellent throughput, otherwise one already needs many channels only to break even. Moreover, one needs to be careful when using gratings because they add temporal dispersion [62]. Furthermore, when computing the resulting expected SNR, one has to pay attention to remaining in the validity range of the equation. In particular, if the light is dispersed over a large number of spectral channels, the optical bandwidth may become very small for each channel, and consequently the coherence time becomes large. With timing resolution on the order of $\sim 20$ ps or even faster [63, 64], the condition $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$ may no longer be fulfilled. It is thus interesting to see how the SNR scaling evolves in that regime. Another critical point is that the spectra we correlate should be the same at the two (or more) telescopes, which becomes more challenging as the spectral channels get narrower. We examine these two points in the following.
2.4.1 Beyond the hypothesis $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$
@@
We consider a Gaussian $g^{(2)}(\tau)$ function [^7] We take a Gaussian instead of a sinc2 (Eq. 12) to avoid inducing a systematic error in the fitting procedure. In a true experiment, the fit function would have to be adapted. of amplitude one and area $\tau_{\mathrm{c}}$, convolved with another Gaussian of area one and rms width $\tau_{\mathrm{el}}$. We add noise as previously and determine the SNR from a 3-parameter Gaussian fit. We report the results in Fig. 6 for three different situations. First, the coherence time is fixed, and we vary $\tau_{\mathrm{el}}$ [Fig. 6(a)]. As expected, it shows that the SNR increases when the timing resolution is improved ($\tau_{\mathrm{el}}$ decreases), but the SNR saturates when $\tau_{\mathrm{el}}$ reaches $\tau_{\mathrm{c}}$. Second, we fix $\tau_{\mathrm{el}}$, and we vary $\tau_{\mathrm{c}}$ by varying the spectral bandwidth $\Delta\nu$ [Fig. 6(b)]. In that case, the number of detected photons increases with $\Delta\nu$, but since $\tau_{\mathrm{c}}$ decreases, in the limit $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$, the SNR becomes independent of the bandwidth, as already discussed in Sec. 2.1. However, in the opposite regime, the (lower) SNR increases as $\sqrt{\Delta\nu}$. This is because the photon flux increases linearly with $\Delta\nu$ and the width of the measured bunching peak is reduced linearly with $\Delta\nu$, as it is fully resolved. Since the fit uncertainty is roughly proportional to the square root of the number of points in the bunching peak, we finally obtain a $\sqrt{\Delta\nu}$ dependence in this regime. Finally, we can also fix the total wavelength range and divide it into $N$ spectral channels, and study how the SNR evolves with $N$ [Fig. 6(c)]. We expect a $\sqrt{N}$ enhancement in the regime $\tau_{\mathrm{c}}\ll\tau_{\mathrm{el}}$, but when $\tau_{\mathrm{c}}>\tau_{\mathrm{el}}$, the SNR saturates. This happens when the number of channels is on the order of the ratio between the total optical bandwidth and the electronic bandwidth. Beyond that point, to continue improving the SNR, one has to increase the total optical bandwidth.
Figure 6: Numerical simulations of the signal-to-noise ratio as a function of $\tau_{\mathrm{el}}$ and $\tau_{\mathrm{c}}$ in different situations. (a) $\tau_{\mathrm{el}}$ is varied and $\tau_{\mathrm{c}}$ is fixed. This corresponds to improving the resolution of the detector (towards the left). The SNR increases as $1/\sqrt{\tau_{\mathrm{el}}}$ as expected until it saturates when $\tau_{\mathrm{el}}\sim\tau_{\mathrm{c}}$. The dashed line is $S\!N\!R=0.44\tau_{\mathrm{c}}/(\sigma_{\mathrm{noise}}\sqrt{\tau_{\mathrm{el}}t_{\mathrm{bin}}})$, with $\tau_{\mathrm{c}}=1$, $t_{\mathrm{bin}}=2\times 10^{-3}$ and $\sigma_{\mathrm{noise}}=10^{-2}$. (b) The detector performance is fixed ($\tau_{\mathrm{el}}=1$) and we vary the optical bandwidth $\Delta\nu$. Although the $x$-axis is the same as in panel (a) in virtue of $\tau_{\mathrm{c}}=1/\Delta\nu$, the result is different because the number of detected photons increases with $\Delta\nu$. In the limit $\Delta\nu\,\tau_{\mathrm{el}}\gg 1$, we recover the well-known result that the SNR is independent of the optical bandwidth. The dashed line is the same equation as in (a) except that now $\sigma_{\mathrm{noise}}=10^{-2}/\Delta\nu$. In the opposite regime, the SNR increases as $\sqrt{\Delta\nu}$. (c) We still consider $\tau_{\mathrm{el}}=1$ and a fixed total wavelength range $\Delta\lambda_{\mathrm{tot}}$ or equivalently a total frequency range $\Delta\nu_{\mathrm{tot}}$, divided into NN independent channels. The computation is the same as in (b) but the SNR is multiplied by $\sqrt{N}$. However, above $N\sim\Delta\nu_{\mathrm{tot}}\,\tau_{\mathrm{el}}$, we enter the regime $\Delta\nu\,\tau_{\mathrm{el}}<1$ and the SNR saturates.
@@
This limitation is implicitly taken into account in the proposal [33], for example, since they consider individual bandwidths of $\Delta\lambda\approx 0.1$ nm at $\lambda\approx 500$ nm with $N\approx 2000$ channels. This corresponds to $\tau_{\mathrm{c}}\approx 8$ ps, very close to the considered timing resolution $\tau_{\mathrm{el}}\approx 10$ ps. But the total wavelength range is then $\Delta\lambda_{\mathrm{tot}}=$ 200 nm, which surely makes the interpretation of the measured visibility a challenge.
@@
Finally, note that there is also a fundamental limit in simultaneously determining the arrival time and frequency (via the spectral dispersion) of a photon [65] : $\Delta\nu\Delta t\geq 1/4\pi$ (Heisenberg or Fourier dispersion relation, where here $\Delta\nu$ and $\Delta t$ are standard deviations). This means that the passage through the dispersive optics induces a supplementary timing uncertainty. The order of magnitude is the same as the coherence time [^8] The exact numerical factors depend on the definitions and spectral shape. For instance, let us consider a Gaussian spectrum of rms width $\Delta\nu$ for simplicity. Then the rms width of the $g^{(2)}(\tau)$ function is ($2\pi\sqrt{2}\Delta\nu)^{-1}$ [39] and the Fourier-limited timing uncertainty $\Delta t=(4\pi\Delta\nu)^{-1}$. The two differ by a factor $\sqrt{2}$.. Therefore, it gives another reason to say that it is useless to improve the timing resolution of detectors below $\tau_{\mathrm{c}}$.
2.4.2 Effect of mismatched spectra